A plane wave is described by an electric field

180πcos10 (2π 100 ).

8 E  t  y

What is its (i) frequency in Hz, (ii) wave number, (iii) phase velocity, 

(iv) wavelength, and (v) direction of propagation?


1
Expert's answer
2021-03-23T08:07:45-0400

Consider a plane wave:


"x(y,t)=180π\text{cos}(10y+2π 100t)."

(i) frequency in Hz:


"f=\frac{\omega}{2\pi}=\frac{2\pi100}{2\pi}=100\text{ Hz}."

(ii) wave number is the term before y:


"k=10\text{ m}^{-1}."

(iii) phase velocity:


"v=\frac{\omega}{k}=\frac{2\pi 100}{10}=20\pi\text{ m/s}."

(iv) wavelength:


"\lambda=\frac{2\pi}{k}=0.2\pi\text{ m}."

(v) direction of propagation: toward positive x.